# Theoretical Modeling of LLM Self-Improvement Training Dynamics Through Solver-Verifier Gap

### v3

@@ C.1 Hyperparameters
C.2 Omitted Figures
In this section, we will provide experiment details including setup and figures omitted in Section 4.
@@ C.2 Omitted Figures
Self Improvement. In this part, we display figures omitted in Section 4. In self-improvement, we employed mini-batch gradient descent with a batch size of $256$. For each model-task pair, we conducted training for $10$ epochs, saving a checkpoint after processing half of the training data. We test the solver accuracy and the verifier accuracy, defined as the accuracy of the response and the BoN response, respectively. Figure 8 illustrates the results of self-improvement on Phi-3.5-mini, Phi-3-mini and Llama-3.2-3B model with QE method. We observe that most of model-task pairs have similar results with the Phi-4-mini model except for several pairs. We also present the results with TF method in Figure 9. When self-improving the Phi-3.5-mini and Phi-3-mini models on Math data set, we observe that accuracy and uncertainty decrease simultaneously after several training epochs. The reason for this phenomenon might be LLM misleading by incorrect responses, as the BoN response may correspond to a incorrect answer with low uncertainty. The solution of this problem could be a future direction. Additionally, we note that the verifier accuracy of Llama-3.2-3B on GSM8k decreases as training progresses and is lower than the solver accuracy after 8 epochs of training. One possible reason for this phenomenon is that we use length-regularized log-likelihood to obtain the BoN responses, so the BoN responses tend to be longer responses. Long responses are more likely to contain repeated content, which may reduce accuracy. In addition, we present the self-improvement results on ProntoQA dataset with TF method Figure 10.
@@
Validation. In this part, we present the validation results of Phi-3.5-mini, Phi-3-mini and Llama-3.2-3B on train split in Figure 11 and observe that the empirical data are strongly fitted to the exponential model for all models. Results of Phi-4-mini, Phi-3.5-mini, Phi-3-mini and Llama-3.2-3B on test split are presented in Figure 12. Experiment results on ProntoQA dataset are presented in Figure 13.
@@
Self Evaluation. This part we provide the results of self-evaluation for different sample size $N$. For each model, we evaluate the accuracy and uncertainty of both the solver and the verifier, varying $N$ across the values 2,4,8,16,32,64{2,4,8,16,32,64}.  Figure 14 illustrates the accuracy and uncertainty of Phi-4-mini, Phi-3.5-mini, Phi-3-mini, Llama-3.2-3B and Llama-3.1-8B on Math and GSM8k with different sample size using TF and QE respectively, which shows that the verifier outperforms solver in all model-task pairs.
@@
Cross Evaluation. In cross evaluation we set $N=16$ because, as demonstrated in Figure 14, the accuracy improve slightly for $N>16$. The results of QE method are presented in Figure 16 while results of TF method are presented in Figure 17. We observe that when a fixed model serves as the solver, the verifier’s performance generally surpasses that of the solver, even when a different model is employed as the verifier.
@@
Pass@K. We investigate the underlying reason for the limit of self-improvement, focusing on how a decrease in the model’s response diversity leads to a diminishing potential energy, thereby causing the model’s capability to plateau. Although the self-improvement paradigm shows great improvement in model capability, it iteratively leads to a performance plateau. To investigate the underlying cause of this saturation, we conduct the Pass@K experiment on the initial model and the self-improved model. Pass@K is a metric that calculates the proportion of prompts where at least one of the $K$ responses is correct. When $K$ is large, Pass@K could be used to measure the diversity of the solver. We present the result in Figure 18. We observe that when $K$ is small, Pass@K increases with the number of epochs of self-improvement, validating the self-improvement process. However, when $K$ is large, we observe a slight decrease in Pass@K, indicating that the diversity of the solver is reduced through self-improvement. The degradation in diversity is caused by the convergence to a certain response, which is a potential reason for the limit of self-improvement. Lack of ProntoQA dataset in experiment as the Pass@K of the QA problem is close to $1$ when $K$ is large.

### v4

@@ D.1 Hyperparameters
D.2 Omitted Figures
In this section, we will provide experiment details including setup and figures omitted in Section˜4.
@@ D.2 Omitted Figures
Self Improvement. In this part, we display figures omitted in Section˜4. In self-improvement, we employed mini-batch gradient descent with a batch size of $256$. For each model-task pair, we conducted training for $10$ epochs, saving a checkpoint after processing half of the training data. We test the solver accuracy and the verifier accuracy, defined as the accuracy of the response and the BoN response, respectively. Figure˜10 illustrates the results of self-improvement on Phi-3.5-mini, Phi-3-mini and Llama-3.2-3B model with QE method. We observe that most of model-task pairs have similar results with the Phi-4-mini model except for several pairs. We also present the results with TF method in Figure˜11. When self-improving the Phi-3.5-mini and Phi-3-mini models on Math data set, we observe that accuracy and uncertainty decrease simultaneously after several training epochs. The reason for this phenomenon might be LLM misleading by incorrect responses, as the BoN response may correspond to a incorrect answer with low uncertainty. The solution of this problem could be a future direction. Additionally, we note that the verifier accuracy of Llama-3.2-3B on GSM8k decreases as training progresses and is lower than the solver accuracy after 8 epochs of training. One possible reason for this phenomenon is that we use length-regularized log-likelihood to obtain the BoN responses, so the BoN responses tend to be longer responses. Long responses are more likely to contain repeated content, which may reduce accuracy. In addition, we present the self-improvement results on ProntoQA dataset with TF method Figure˜12.
@@
Validation. In this part, we present the validation results of Phi-3.5-mini, Phi-3-mini and Llama-3.2-3B on train split in Figure˜14 and observe that the empirical data are strongly fitted to the exponential model for all models. Results of Phi-4-mini, Phi-3.5-mini, Phi-3-mini and Llama-3.2-3B on test split are presented in Figure˜15. Experiment results on ProntoQA dataset are presented in Figure˜16.
@@
Self Evaluation. This part we provide the results of self-evaluation for different sample size $N$. For each model, we evaluate the accuracy and uncertainty of both the solver and the verifier, varying $N$ across the values 2,4,8,16,32,64{2,4,8,16,32,64}.  Figure˜18 illustrates the accuracy and uncertainty of Phi-4-mini, Phi-3.5-mini, Phi-3-mini, Llama-3.2-3B and Llama-3.1-8B on Math and GSM8k with different sample size using TF and QE respectively, which shows that the verifier outperforms solver in all model-task pairs.
@@
Cross Evaluation. In cross evaluation we set $N=16$ because, as demonstrated in Figure˜18, the accuracy improve slightly for $N>16$. The results of QE method are presented in Figure˜21 while results of TF method are presented in Figure˜22. We observe that when a fixed model serves as the solver, the verifier’s performance generally surpasses that of the solver, even when a different model is employed as the verifier.
@@
Pass@K. We investigate the underlying reason for the limit of self-improvement, focusing on how a decrease in the model’s response diversity leads to a diminishing potential energy, thereby causing the model’s capability to plateau. Although the self-improvement paradigm shows great improvement in model capability, it iteratively leads to a performance plateau. To investigate the underlying cause of this saturation, we conduct the Pass@K experiment on the initial model and the self-improved model. Pass@K is a metric that calculates the proportion of prompts where at least one of the $K$ responses is correct. When $K$ is large, Pass@K could be used to measure the diversity of the solver. We present the result in Figure˜23. We observe that when $K$ is small, Pass@K increases with the number of epochs of self-improvement, validating the self-improvement process. However, when $K$ is large, we observe a slight decrease in Pass@K, indicating that the diversity of the solver is reduced through self-improvement. The degradation in diversity is caused by the convergence to a certain response, which is a potential reason for the limit of self-improvement. Lack of ProntoQA dataset in experiment as the Pass@K of the QA problem is close to $1$ when $K$ is large.
@@
Figure 9: Self-improvement dynamics across Base, Mid-trained, and Post-trained stages on test split.
@@
Figure 13: Accuracy and uncertainty during the self-improvement of the Phi-4-mini, Phi-3.5-mini, Phi-3-mini and Llama-3.2-3B on MBPP datset using the TF method.
@@
(a) Train Split
(b) Test Split
Figure 17: Uncertainty curve versus epochs of Phi-4-mini, Phi-3.5-mini, Phi-3-mini and Llama-3.2-3B models using TF method on MBPP datasets.
@@
Figure 20: Accuracy and uncertainty of Phi-4-mini, Phi-3.5-mini, Phi-3-mini and Llama-3.2-3B on MBPP dataset across different sample size using TF method.
@@
Appendix E Additional Discussion
Phenomenological results. We acknowledge that neural network optimization is complex. However, our approach is phenomenological, akin to how thermodynamics describes complex molecular systems using macroscopic variables. Our core assumption, the linear approximation of potential energy is empirically validated. As shown in Figure $2$ and Figure $7$, the relationship between the gap $G$ and its rate of change $dG/dt$ is strictly linear, with $R^{2}>0.9$. Based on linear assumption, exponential convergence law (Eq 10 & Eq 11) fits the training trajectories shown in Figure 4, with $R^{2}>0.9$. Although deriving this relationship from first principles remains an open challenge, our framework yields valuable insights into the self-improvement mechanism, it provides a precise mathematical characterization of the optimization curve and a rigorous basis for modeling cross-improvement strategies.
Simplified assumptions (linear and time-invariant). We clarify that linearity is inspired by empirical observations, not just a prior assumption, and the time-invariance is a local approximation. Figures 2 and 7 plot the $dG/dt$ against $G(t)$, revealing a linear relationship between the potential energy $E(t)$ and gap $G(t)$ across various models.
  • Model Complexity and Parameter Estimation. Modeling the coefficients ($\alpha,\beta$) as time-varying functions would significantly increase the complexity of the theoretical framework. Introducing time-varying parameters would result in complex modeling processes and make parameter estimation nearly impossible without introducing further arbitrary assumptions. Given that our constant-coefficient model already achieves a high degree of precision (with $R^{2}>0.9$ in fitting training dynamics), suggesting the time-invariant assumption is the most effective approach for capturing the macroscopic dynamics.
  • Time-Homogeneity of the State. In our framework, the dynamics are driven by the state of the system (the Solver-Verifier Gap) rather than explicit time. This implies time-homogeneity: any time point $t$ in the training process can be regarded as a start point. The evolution of the system depends on the current capability gap $G(t)$ and the potential energy $E(t)$. This is empirically supported by Figure 2, which shows a consistent linear relationship between the gap and its rate of change throughout the training process, confirming that the parameters governing this relationship remain stable.
  • Applicability of Time-Varying Parameters. We acknowledge that the time-invariant assumption is a local approximation valid for the standard self-improvement regimes studied in this paper. In specific complex scenarios, such as training across radically different data distributions or using highly dynamic learning rate schedules, the coefficients might drift. We have identified relaxing the time-invariant property as a specific direction (e.g., piece-wise constant) for future work to extend the framework’s generality.
Connection between theoretical variables and practical LLM learning signals. The variables in our phenomenological model map directly to concrete, measurable quantities in the LLM training process:
  • The capability gap $G(t)$: As defined in $Eq.(6),G(t)=U_{v}(t)-U_{s}(t)$. This represents the "Quality Delta" between the model’s current generation performance $U_{s}$, and the quality of the verification output $U_{v}$. In practice, this is measured by comparing the average NLL of the model’s outputs against the average NLL of the Best-of-N responses selected for training.
  • The potential energy $E(t)$: E(t) is a function of the gap, approximated linearly as E(t)=kG(t)bE(t)=kG(t)-b. This represents the power which drives self-improvement. Equation $8$ states $\frac{dU_{s}(t)}{dt}=-\alpha E(t)$. This means $E(t)$ maps directly to the speed of improvement of the model. In a practical training run, a high $E(t)$ indicates that the gap is large enough to provide strong gradients, leading to a rapid drop in loss. A low $E(t)$ indicates the model has exhausted the information in the current gap, leading to a training plateau.
Therefore, monitoring $E(t)$ and $G(t)$ might be equivalent to monitoring the efficiency of the data relative to the model’s current state.
Validations of cross-improvement (Equation 12). We justify its validity from both theoretical and empirical perspectives:
  • Theoretical Rationality: Equation 12 models the capability enhancement as a gain factor mechanism. The term $(1+\gamma\eta_{t})^{-1}$ acts as a discount factor on the uncertainty. As the ratio of high-quality external data $\eta_{t}$ increases, the verifier’s uncertainty $U_{v}$ decreases monotonically. This phenomenological formula captures the intuition that external data reduce the system’s entropy, with $\gamma$ representing the efficiency of this external information. From a technical standpoint, we explicitly chose this functional form for its mathematical tractability. This specific inverse-linear form allows the coupled dynamics to be approximated by linear matrix operations. If we were to use more complex non-linear scaling laws, the system would become analytically intractable, preventing the derivation of the closed-form approximate solution in Proposition 5.1. This form strikes a necessary balance between capturing physical intuition and maintaining solvability.
  • Empirical Validation: Crucially, the validity of this modeling assumption is supported by our experimental results. Based on Eq.12, we derived Proposition 5.1, which predicts that the final solver capability depends primarily on the total amount of external data, rather than its specific allocation timing. Our experiments in Section 5.2 (Table 1) confirm this conclusion. We observed that different allocation strategies (Early, Uniform, Late) yielded negligible differences in final accuracy. This alignment between the theoretical prediction derived from Eq.12 and the empirical outcome strongly validates the reasonableness of the model.
Results and difference of model families. Different model families exhibit distinct exponential trajectories because their inherent properties (such as pre-training data quality and architectural maturity) determine the specific values of the decay rate coefficient and the boundary conditions in our derived analytical solution. As discussed in Q1, Llama families exhibit a larger solver learning rate $\alpha$. Mathematically, a larger $\alpha$ maximizes the exponent magnitude $|-k(\alpha-\beta)|$, resulting in a steeper decay curve (rapid convergence), as observed in our Llama experiments. Phi families are pre-trained on high-quality synthetic data. This results in a smaller gap and decay constant, leading to a flatter decay function. Crucially, the decay function is mathematically coupled to the limit. The theoretical asymptote is defined as $U_{s,\infty}=U_{s,0}-\frac{\alpha}{\alpha-\beta}(G_{0}-\frac{b}{k})$.
  • Phi Family: Starts with a significantly smaller Initial Gap ($G_{0}$) and lower initial uncertainty ($U_{s,0}$) due to high-quality upstream synthesis. Plugging these smaller values into the equation mathematically dictates a better final capability.
  • Llama Family: Starts with a larger $G_{0}$ and $U_{s,0}$. Despite the rapid initial descent, the higher starting baseline mathematically leads to convergence at a relatively higher asymptote.
Potential energy as driven force. We justify modeling the training dynamics as Equation˜9:
  • Physical Analogy: Potential Difference as the Driving Force. In this paper, we model the training dynamics as a form of $\frac{dU_{s}}{dt}=-\alpha E(t),\frac{dU_{v}}{dt}=-\beta E(t)$, which can be analogous to Newton’s Second Law. In Newton’s Second Law ($F=ma$), force is the cause that generates acceleration. Without force, there is no change in the state of motion. Similarly, in our framework, the Potential Energy $E(t)$ acts as the force which is the cause that drives the evolution of the model’s capabilities ($U_{s},U_{v}$)(inspired by Song et al. (2025)). Just as a physical force dictates how an object moves, the magnitude of this potential energy dictates the intensity of the self-improvement process. This modeling captures the core intuition: the existence of a gap creates a potential difference that physically necessitates and drives the optimization process.
  • Empirical Validation. Despite the simplicity of our model, our results demonstrate that this simple macroscopic model captures the training dynamics with remarkable precision. We fit the theoretical exponential curves derived from our conservative field model to the actual training trajectories. As shown in Figure 4, despite the model’s simplicity, it fits the empirical training data (both Math and GSM8k datasets) exceptionally well, with coefficients of determination ($R^{2}$) consistently exceeding $0.9$.
Conservative vector field. To statistically verify whether historical information (G(t1)G(t-1)) provides any necessary explanatory power for the dynamics, we performed a Partial F-test to compare two nested regression models:
  • Model 1 (Simple Model): $\frac{dG}{dt}=\beta_{1}G(t)+\epsilon$
  • Model 2 (Full Model): $\frac{dG}{dt}=\beta_{1}G(t)+\beta_{2}G(t-1)+\epsilon$
We evaluated whether adding the historical term G(t1)G(t-1) significantly reduces the Residual Sum of Squares (RSS). As shown in Table˜4, the insignificant F-test result leads us to fail to reject the null hypothesis ($\beta_{2}=0$). This statistically proves that adding historical information does not significantly improve the model’s fit. The dynamics are adequately described by the current state alone, providing robust evidence for the memoryless nature of the potential energy.
Table 4: Model Selection via Partial F-test.
Model Dataset Split F-statistic $P$-value (F)
Phi-4-mini Math Train 1.295 0.2687
Math Test 1.203 0.2911
GSM8k Train 0.024 0.8771
GSM8k Test 9.277 0.0122
Llama-3.2-3B Math Train 0.479 0.5114
Math Test 0.833 0.3917
GSM8k Train 0.875 0.3808
GSM8k Test 0.808 0.3985
Phi-3.5-mini Math Train 1.807 0.2209
Math Test 3.218 0.1159
GSM8k Train 1.945 0.2058
GSM8k Test 3.140 0.1197
Phi-3-mini Math Train 0.000 0.9845
Math Test 5.300 0.0548
GSM8k Train 0.528 0.4911
GSM8k Test 2.163 0.1848